OpenStudy (anonymous):

Aperfectly square lot has a perimeter of exactly 1 mile. A. find the length of one side B. Find the area

jimthompson5910 (jim_thompson5910):

A) Length of one side = Perimeter/4 = 1/4 So the length of one side is 1/4 mile (ie a quarter mile) B) Area = side * side = (1/4)*(1/4) = 1/16 So the area is 1/16 square miles (ie one-sixteenth of a square mile)

OpenStudy (anonymous):

so the answer for length is 1/4 and the area is 1/16

jimthompson5910 (jim_thompson5910):

yep, don't forget the units

OpenStudy (anonymous):

okayy thanksss